OpenStudy (anonymous):

Find the equation of the line that contains the given point and has the given slope. Let y be the dependent variable and let x be the independent variable. Point (-1, -1), m =1/7 This is what I think you suppose to do but its not working for me.... y-y1=m(x-x1) y=1/7x+1? am i doing this right?

OpenStudy (anonymous):

y=1/7x+1/7?

OpenStudy (anonymous):

y+1=1/7(x+1) y+1=1/7x+1/7 y=1/7x-6/7

OpenStudy (anonymous):

where did 6/7 come from?

OpenStudy (anonymous):

i subtracted 1 from each side

OpenStudy (anonymous):

Im sorry i confused is there a way to show me how you did it? u subtracted what from both sides?

OpenStudy (anonymous):

point-slope formula is the top line y-y1=m(x-x1) distributed m on the 2nd line y-y1=mx-mx1 added y1 to each side on 3rd line y=mx-mx1+y1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

in your solution, i believe you forgot to distribute the 1/7 to the -1

OpenStudy (anonymous):

to the positive 1, rather

OpenStudy (anonymous):

I did and I see what you did now that was just my missing step. Thank you so much

OpenStudy (anonymous):

np